OpenStudy (anonymous):

How do I solve this system of equations? y=x^2+3 y=x+5

OpenStudy (firekat97):

you have both equations equalling y, so why not let both of them equal one another? try that and simplify

OpenStudy (firekat97):

so to get you started, you should have something like this... x^2 + 3 = x + 5 so x^2 - x - 2= 0 do you understand why?

OpenStudy (anonymous):

Yep I got X=2 or -1 as my solutions

OpenStudy (firekat97):

Really sorry for the super late response! anyway, yup thats correct! now that you have your have your x solutions out, substitute them into either equations to get out your y solutions, and there you have it. :)